Please solve this problem: Raisins cost $4.00/lb and peanuts cost $3.00/lb. A 60lb mixture or raisins and peanuts cost $3.40/lb. How many pounds of each type of food is in the mixture?
Answer by scott8148(6628)

"A 60lb mixture or raisins and peanuts" ___ r + p = 60 ___ r = 60 - p


($3.00)(p) + ($4.00)(60 - p) = (3.40)(60)